# How Transformers Solve Propositional Logic Problems: A Mechanistic Analysis

### v3

@@ 1.1 Related works
Evaluation of reasoning abilities of LLMs. Our work is also related to the line of work which focus on empirically evaluating the reasoning abilities of LLMs across different types of tasks (Xue et al., 2024; Tafjord et al., 2021; Hendrycks et al., 2021; Chen et al., 2024; Patel et al., 2024; Berglund et al., 2024; Morishita et al., 2023; Fu et al., 2024; Seals & Shalin, 2024; Zhang et al., 2023; Saparov & He, 2023; Saparov et al., 2024; Luo et al., 2024; Arora et al., 2024; Han et al., 2024; Dziri et al., 2024; Yang et al., 2024). While these studies primarily benchmark their performance on sophisticated tasks, our work focuses on understanding “how” transformers reason on logic problems accessible to fine-grained analysis.
Analysis of how LLMs reason. There are far fewer studies that focus on providing fine-grained analysis of how LLMs reason. To the best of our knowledge, only a handful of works, such as Wang et al. (2023); Zečević et al. (2023); Ye et al. (2024), share similar goals of understanding how transformers perform multi-step reasoning through detailed empirical or theoretical analysis. However, none studies the [Variable relationships]+[Variable value assignment]+[Query] type problem in conjunction with analysis on both small transformers trained purely on the synthetic problem, and large language models trained on a large corpus of internet data. Moreover, there is little study on the existence and properties of circuits employed by LLMs (in the wild) for performing latent in-context multi-hop reasoning.
@@
2 Data model: a propositional logic problem
Figure 1: Synthetic data model. The causal structure has two chains: one with a logical operator (LogOp) at the end and the other being purely a linear causal chain. They do not share any proposition variable. This example is the length-3 case and all symbols are explained in §2.
In this section, we describe the synthetic propositional logic problem that shall be the data model of this paper. Our problem follows an implicit causal structure, as illustrated in Figure 1. The structure consists of two distinct chains: exactly one containing a logical operator at the end of the chain, and one forming a purely linear chain. The two chains do not share any proposition variable.
We require the model to generate a minimal reasoning chain, consisting of “relevant facts”, proper rule invocations, and intermediate truth values, to answer the truth-value query. Consider an example constructed from the causal graph in Figure 1, written in English:
  • Rules: K implies D. D or E implies A. V implies E. T implies S. P implies T.
  • Facts: K is true. P is true. V is false.
  • Query: A.
  • Answer: K is true. K implies D; D is true. D or E implies A; A is true.
In this example, the QUERY token A is the terminating node of the OR chain. Since any true input to an OR gate (either D or E) results in A being true, the minimal solution chooses only one of the starting nodes from the OR chain to construct its argument: in this case, node K is chosen.
Minimal proof and solution strategy. In general, the problem requires a careful examination of the rules, facts and query to correctly answer the question. First, the QUERY token determines the chain to deduce its truth value. Second, if it is the logical-operator (LogOp) chain being queried, the model needs to check the facts to determine the correct facts to write down at the start of the reasoning steps (this step can be skipped for queries on the linear chain). Third, the proof requires invoking the rules to properly deduce the truth value of the query token.
3 The reasoning circuit in a small transformer
  1. 1.
  2. 2.
3.1 Learner: a decoder-only attention-only transformer
In this section, we study decoder-only attention-only transformers, closely resembling the form of GPT-2 (Radford et al., 2019a). We train these models exclusively on the synthetic logic problem. The LogOp chain is queried 80% of the time, while the linear chain is queried 20% of the time during training. Details of the model architecture are provided in Appendix B.2.
Architecture choice for mechanistic analysis. We select a 3-layer 3-head transformer to initiate our analysis since it is the smallest transformer that can achieve 100% test accuracy; we also show the accuracies of several candidate model sizes in Figure 6 in Appendix B for more evidence. Note that a model’s answer on a problem is considered accurate only if every token in its answer matches that of the correct answer. Please refer to Appendix B.2 for an illustration of the model components.
3.2 Mechanism analysis
  1. 1.
  2. 2.
  3. 3.
We illustrate the overall reasoning strategy and core evidence for it in Figure 8 in Appendix B.3.
3.2.1 Linear or LogOp chain: routing signal at the QUERY position
Routing direction at QUERY. There exists a “routing” direction ${\bm{h}}_{route}$ present in the embedding generated by the layer-2 attention block, satisfying the following properties:
  1. nosep
    $\alpha_{1}({\bm{X}}){\bm{h}}_{route}$ is present in the embedding when the linear chain is queried, and $\alpha_{2}({\bm{X}}){\bm{h}}_{route}$ is present when the LogOp chain is queried, where the two $\alpha_{i}({\bm{X}})$’s are sample dependent, and satisfy the property that $\alpha_{1}({\bm{X}})>0$, and $\alpha_{2}({\bm{X}})<0$.
  2. nosep
    The “sign” of the ${\bm{h}}_{route}$ signal determines the “mode” which layer-3 attention operates in at the ANSWER position. When a sufficiently “positive” ${\bm{h}}_{route}$ is present, layer-3 attention acts as if QUERY is for the linear chain by placing significant attention weight at the QUERY position. A sufficiently “negative” ${\bm{h}}_{route}$ causes layer-3 to behave as if the input is the LogOp chain: the model focuses attention on the rules and fact sections, and in fact outputs the correct first token of the LogOp chain!
Figure 2: Cosine similarity matrix between output embeddings from layer-2 attention block. Samples 0 to 99 query for the linear chain, samples 100 to 199 query for the LogOp chain. Observe the in-group clustering in angle (top left and bottom right), and the negative cross-group cosine similarity (top right and bottom left).
Evidence 1b: distinct layer-3 attention behavior w.r.t. chain type. When the linear chain is queried, the layer-3 attention heads predominantly focus on the QUERY position, with over 90% of their attention weights on the QUERY position on average (based on 1k test samples). In contrast, when the LogOp chain is queried, less than 5% of layer-3 attention is on the QUERY on average. Instead, attention shifts to the Rules and Facts sections of the context, as shown in Figure 9 in Appendix B.4.
Evidence 1c: computing ${\bm{h}}_{route}$, and proving its role with interventions. To erase the instance-dependent information, we average the output of the second attention block over 1k samples where QUERY is for the linear chain. We denote this estimated average as $\hat{{\bm{h}}}_{route}$ which effectively preserves the sample-invariant signal. To test the influence of $\hat{{\bm{h}}}_{route}$, we investigate its impact on the model’s reasoning process, and we observe two intriguing properties:
  1. 1.
[…]
  2. 2.
[…]
It follows that ${\bm{h}}_{route}$ indeed exists, and the “sign” of it determines the attention patterns in layer 3 (and the overall network’s output!) in the aforementioned manner.
3.2.2 Answer for the linear and LogOp chain
  1. 1.
  2. 2.
We delay the full set of evidence for the above two claims to Appendix B.4 due to space limits. Our argument mainly relies on linear probing and causal interventions at different layers and token positions in the model.
4 The reasoning circuit in LLMs
We now turn to examine how pretrained LLMs, namely Mistral-7B-v0.1 and Gemma-2-9B, solves this reasoning problem. We choose these LLMs as they are amongst the smallest accessible models which achieves above 70% accuracy on (a minimal version of) our problem. Our primary focus here is the same as in the previous section: how does the model infer the first answer token without any CoT? We are interested in this question as the first answer token requires to model to process all the information in the context properly without access to any CoT, and demands the greatest number of latent reasoning steps out of all answer tokens.
Figure 3: High-level properties of Mistral-7B’s reasoning circuit. The (chunks of) input tokens are on the left, which are passed into the residual stream and processed by the attention heads. We illustrate the information flow manipulated by the different types of attention heads we identified to be vital to the reasoning task.
In the main text, we primarily focus on presenting the reasoning circuit discovery and verification results of Mistral-7B, and delay the results for Gemma-2-9B to Appendix C.7. We illustrate the reasoning circuit inside Mistral-7B for this prediction task in Figure 3. At a high level, there are two intriguing properties of the reasoning circuit of the LLM: [^1] We use $(\ell,h)$ to denote an attention head. When referencing multiple heads in the same layer, we write $(\ell,h_{1};h_{2};...;h_{n})$ for brevity.
  1. 1.
  2. 2.
    There is a sparse set of attention heads that are found to be central to the reasoning circuit: the queried-rule locating heads, queried-rule mover heads, fact-processing heads, and decision heads. We discuss circuit discovery in §4.3.1, and circuit verification in §4.3.2. We summarize the high-level conclusions of our more involved experiments in §4.3.3, and refer the reader to the respective sections in the Appendix from there.
4.1 Minimal problem description
In our LLM experiments, the input samples have the following properties:
  1. 1.
    Before asking for the answer, we give Mistral-7B 6 randomly chosen in-context examples, and 4 for Gemma-2-9B.
  2. 2.
    The problem is length-2: only one rule involving the OR gate, and one linear-chain rule. Moreover, the answer is always true. In particular, the truth values of the two premise nodes of the OR chain always have one FALSE and one TRUE.
  3. 3.
    The proposition variables are all (single-token) capital English letters.
The design decision in the first point is to ensure fairness to the LLM which was not trained on our specific logic problem. As for the last two points, we restrict the problem in this fashion mainly to ensure that the first answer token is unique, which improves the tractability of the analysis. Note that these restrictions do not take away the core challenge of this problem: the LLMs still need to process all the context information without CoT to determine the correct first answer token, and out of all answer token positions, the first token still requires the greatest number of latent reasoning steps to resolve. We discuss concrete examples and finer details of testing the models on this problem in Appendix C.
4.2 Causal mediation analysis
We provide evidence in this part of the paper primarily relying on a popular technique in mechanistic interpretability: causal mediation analysis. Our methodology is as follows:
  1. 1.
    Suppose we are interested in the role of the activations of certain components of the LLM in a certain (sub-)task. For a running example, say we want to understand what role the attention heads play in processing and passing QUERY information to the “:” position for inference. Let us denote the activations as ${\bm{A}}_{\ell,h;t}({\bm{X}})$, representing the activation of head $h$ in layer $\ell$, at token position $t$.
  2. 2.
    Typically, the analysis begins by constructing two sets of prompts which differ in subtle ways. A natural construction in our example is as follows: define sets of samples $\mathcal{D}_{orig}$ and $\mathcal{D}_{alt}$, where ${\bm{X}}_{orig,n}$ and ${\bm{X}}_{alt,n}$ have exactly the same context, except in ${\bm{X}}_{orig,n}$, QUERY is for the LogOp chain, while in ${\bm{X}}_{alt,n}$, QUERY is for the linear chain. Moreover, denote the correct targets $y_{orig,n}$ and $y_{alt,n}$ respectively.
  3. 3.
    We run the LLM on $\mathcal{D}_{orig}$ and $\mathcal{D}_{alt}$, caching the attention-head activations. We also obtain the logits of the model. We can compute the model’s logit differences
    $\Delta_{orig,n}=\text{logit}({\bm{X}}_{orig,n})[y_{orig,n}]-\text{logit}({\bm{X}}_{orig,n})[y_{alt,n}],$ (1)
    $\Delta_{alt,n}=\text{logit}({\bm{X}}_{alt,n})[y_{alt,n}]-\text{logit}({\bm{X}}_{alt,n})[y_{orig,n}].$
    For a high-accuracy model, $\Delta_{orig,n}$ and $\Delta_{alt,n}$ should be large and positive for most $n$’s, since most of the time it must be able to clearly tell that on an ${\bm{X}}_{orig,n}$, it is the LogOp chain which is being queried, not the linear chain (similarly on an ${\bm{X}}_{alt,n}$).
  4. 4.
    We now perform intervention for all $n,\ell,h$ and $t$:
    1. nosep
      Run the model on ${\bm{X}}_{orig,n}$, however, replacing the original activation ${\bm{A}}_{\ell,h;t}({\bm{X}}_{orig,n})$ by the altered ${\bm{A}}_{\ell,h;t}({\bm{X}}_{alt,n})$. Now let the rest of the run continue. [^2] Please note that layers $\ell+1$ to $L$ are influenced at and after token position $t$, and technically speaking, now operate “out of distribution”. Let us denote the logits obtained in this intervened run as $\text{logit}^{\to alt;(\ell,h,t)}({\bm{X}}_{orig,n})$.
    2. nosep
      Now compute the intervened logit difference
      $\Delta_{orig\to alt,n;(\ell,h,t)}=\text{logit}^{\to alt;(\ell,h,t)}({\bm{X}}_{orig,n})[y_{alt,n}]-\text{logit}^{\to alt;(\ell,h,t)}({\bm{X}}_{orig,n})[y_{orig,n}].$
  5. 5.
    Average the $\Delta_{orig\to alt,n;(\ell,h,t)}$’s over $n$ for every $\ell,h$ and $t$ (recall that $n$ is the sample index).
  6. 6.
    This procedure helps us identify components that are significant in processing the QUERY information for inference. Intuitively, an activation that result in a positive and large $\Delta_{orig\to alt,n;(\ell,t)}$ (with an ideal upper limit being $\Delta_{alt,n}$) play a significant role in this subtask, because this activation helps “altering” the model’s “belief” from “QUERY is for the LogOp chain” to “QUERY is for the linear chain”. We denote the components identified as a set $\mathcal{C}$.
  7. 7.
    Remark: due to the symmetry of this running example, it is perfectly sensible to perform $alt\to orig$ interventions too, by mirroring the above procedures; we indeed adopt this mirrored procedure too in our experiments for efficient use of model computations.
QUERY-based activation patching is important to reasoning circuit discovery and verification. To justify this claim, there are two points to emphasize first. (1) To solve the reasoning problem, the QUERY is critical to initiating the reasoning chain: without it, the rules and facts are completely useless; with it, the reasoner can then proceed to identify the relevant rules and facts to predict the answer. (2) The prompt pairs differ only by the QUERY token: there is complete information for the model to provide answer for both chains given the QUERY.
If performing the aforementioned interventions on a model component leads to a large intervened logit difference (i.e. it alters the model’s “belief”), then this component must be integral to the reasoning circuit, because the component is now identified to be QUERY-sensitive and has causal influence on (parts of) the model’s reasoning actions. Furthermore, if on (most of) the prompt pairs, by patching all the circuit components in $\mathcal{C}$ simultaneously while freezing the rest causes the intervened logit difference to approach the “maximal” altered logit difference [^3] In this case, we simply consider the “maximal” altered logit difference to be the logit difference of the model run on the altered prompts without interventions, i.e. the $\Delta_{alt,n}$’s., then we obtain evidence suggesting sufficiency of $\mathcal{C}$ for explaining the LLM’s (QUERY-sensitive) reasoning actions.
For greater clarity, we provide illustrations of the intervention experiments in Appendix C.2.
4.3 Circuit analysis
In this section, we discuss properties of the reasoning circuit of Mistral-7B. The order by which we present the results will roughly follow the process which we discovered and verified the circuit; we believe this adds greater transparency to the circuit analysis process. We delay the more involved (and complete) set of experimental results to Appendix C.
4.3.1 Circuit discovery: necessity-based search for model components
We initiate our analysis with QUERY-based patching, following the same procedure as detailed in §4.2. In this set of experiments, we discover the main attention heads responsible for processing the context and performing inference as introduced in the beginning of this Section.
High-level interventions. Figure 4(a) helps us locate a small set of attention heads which are central to the “belief altering” of the LLM. More specifically, only attention heads (12,9), (13,11;22), (14,24;26), (16,0;12;14), (17,25), (19,8;9;16), and (9,25;26) [^4] We discover these two heads with a slightly different patching experiment, presented in Appendix C.3.2. are observed with relatively high intervened logit differences. As for the MLPs, shown in Figure 4(b), play little role in this circuit, except for MLP-0. However, MLP-0 had been observed to act more as a “nonlinear token embedding” than a complex high-level processing unit (Wang et al., 2023). In the rest of this section, we primarily devote our analysis to the attention heads, and leave the exact role of the MLPs to future work.
Figure 4: Attention head and MLP output patching results (over all token positions in relevant context) in (a) and (b) respectively. The closer to 1, the greater the causal influence a component has on the LLM’s correct inference. We highlight the heads with the highest intervened logit difference in (a). We observe that only a small set of attention heads, and the first MLP layer exhibit strong causal influence on the model’s correct inference.
Remark. In Figure 4 and the rest, unless otherwise specified, we adopt a calibrated version of the logit difference (see Appendix C): the closer to 1, the more significant the component is in “altering” the “belief” of the model on the selected subtask.
Attention-head sub-component patching (QUERY-based patching). We now aim to understand why the attention heads identified in the last sub-section are important. For now, we continue with QUERY altering in the prompt pairs. Through intervening on the sub-components of each attention head, namely their value, key, and query, and through examining details of their attention weights, we find that there are roughly four types of attention heads. We show the results in Figure 5.
  1. nosep
    Queried-rule locating head. Attention head (12,9)’s query activation has a large intervened logit difference according to Figure 5(a), therefore, its query and attention patterns are QUERY-dependent and contribute to altering the model’s “belief”. Furthermore, at the QUERY position, we find that on average, its attention weight is above 90% at the “conclusion” variable of the rule being queried. In other words, it is responsible for locating the queried rule, and storing that rule’s information at the QUERY position.555(9,25;26), (14,24;26) exhibit similar tendencies, albeit with smaller intervened logit differences.
  2. nosep
    Queried-rule mover head. Attention head (13,11)’s value activations have large intervened logit difference, and intriguingly, its query and key activations do not share that tendency. This already suggests that its attention pattern performs a fixed action on both the original and altered prompts, and only the value information is sensitive to QUERY. Furthermore, within the relevant context (excluding the 6 in-context examples given), (13,11) assigns above 50% attention weight to the QUERY position, and its attention weight at QUERY is about 10 times larger than the second largest one on average. Recalling the role of layer 12, we find evidence that (13,11) moves the QUERY and queried-rule information to the “:” position.666(13,22), (15,8), (16,0) also appear to belong to this type, albeit with smaller intervened logit difference.
  3. nosep
    Fact processing heads. Attention heads (16,12;14) and (14,26)’s query activations have large intervened logit differences. Within the relevant context, at the “:” token position, their attention weights are above 56, 80 and 63% respectively in the fact section of the context (starting from “Fact” and ending on “.” before “Question”).
  4. nosep
    Decision head. Attention head (19,8)’s query activations have large intervened logit differences. Its attention pattern suggests that it is a “decision” head: within the relevant context, when the model is correct, the head’s top-2 attention weights are always on the correct starting node of the queried rule and the correct variable in the fact section, and the two token positions occupy more than 60% of its total attention in the relevant context on average. In other words, it already has the answer.777(17,25), (19,9;16) exhibit similar tendencies, albeit with smaller intervened logit differences.
Figure 5: Patching of query and value activations in (a) and (c); we found that intervening the key activations only yield trivial scores, so we do not report them here. We show in (b) the typical attention patterns of a representative set of the attention heads which are identified to be important in the intervention experiments shown in (a) and (c). There are several distinct observations which can be made in (b). Queried-rule locating head (12,9): observe that it correctly locates the queried rule which ends with Q. Queried-rule mover head (13,11): the only token position which it focuses on is the QUERY token Q. Fact processing head (16,14): attention concentrates in the fact section. Decision head (19,8): attention focused on the correct first answer token K.
We delay detailed inspection and visualization of the attention statistics to Appendix C.3.3.
4.3.2 Circuit verification: a sufficiency test
A natural question now arises: is $\mathcal{C}$ sufficient to explain the (QUERY-sensitive) reasoning actions of the LLM? As explained in §4.2, we quantify sufficiency by measuring the “belief” altering effect of the circuit, and test $\mathcal{C}$ as follows. Given the original and altered prompt pairs $\{({\bm{X}}_{orig,n},{\bm{X}}_{alt,n})\}_{n\in[N]}$ which differ only by QUERY, we aim to verify that, simultaneously patching ${\bm{A}}_{\ell,h;t}({\bm{X}}_{orig,n})\to{\bm{A}}_{\ell,h;t}({\bm{X}}_{alt,n})$ for every attention head in the circuit $\mathcal{C}=\{(12,9),(13,11),(16,12),(19,8),...\}$ at their functioning token positions, while freezing all the other attention heads to the original activations ${\bm{A}}_{\ell,h;t}({\bm{X}}_{orig,n})$, lead to the average circuit-intervened logit difference $\Delta^{\mathcal{C}}_{orig\to alt}=\frac{1}{N}\sum_{n=1}^{N}\Delta_{orig\to alt,n}^{\mathcal{C}}$ approaching or surpassing the average logit difference of the (un-intervened) model run on the altered prompts, namely $\Delta_{alt}=\frac{1}{N}\sum_{n=1}^{N}\Delta_{alt,n}$. We confirm this hypothesis in Table 1 below.
$\Delta^{\mathcal{C}^{\dagger}}_{orig\to alt}/\Delta_{alt}$1.0 0.98 -1.02 -0.99 -0.25 -0.89
Table 1: $\Delta^{\mathcal{C}^{\dagger}}_{orig\to alt}/\Delta_{alt}$, with different choices of $\mathcal{C}^{\dagger}$. $\mathcal{C}_{null}$ denotes the empty circuit, i.e. the case where no intervention is performed. We abbreviate the attention head families, for example, $DH=$ decision heads; $\mathcal{C}-DH=$ full circuit but with the decision heads removed.
To be precise, the circuit which we perform verification on is the union of the four attention head families, $\mathcal{C}=QRLH\cup QRMH\cup FPH\cup DH$ on their respective functioning token positions:
[…]
We find that by patching all 14 attention heads in $\mathcal{C}$, $\Delta^{\mathcal{C}}_{orig\to alt}$ is about 98% of the “maximal” average logit difference $\Delta_{alt}$ on the altered samples. Moreover, removing any one of the four families of attention heads from $\mathcal{C}$ in the circuit interventions renders the “belief altering” effect of the intervention almost trivial. We present further discussions of the experimental procedure and results, and caveats of reasoning circuit verification in Appendix C.4.
4.3.3 Deeper examinations
In this subsection, we summarize the more involved experiments we performed in addition to the ones we presented so far, and refer the reader to their respective sections in the Appendix for further reading.
Early evidence of a similar reasoning circuit in Gemma-2-9B (§C.7). We performed a similar set of circuit discovery and verification experiments on Gemma-2-9B, and found that the set of attention heads in its circuit can also be divided into four families, closely resembling the attention patterns of the four families in Mistral-7B. In particular, Gemma-2-9B’s circuit is formed by the following families of attention heads, with $\mathcal{C}_{gemma}=QRLH\cup QRMH\cup FPH\cup DH$:
[…]
  • $DH=\text{Decision Heads}=\{(28,12),(30,9)\}$.
[…]
Further validating evidence of the attention head families in Mistral-7B (§C.5, C.6). We present further validating evidence of the queried-rule locating heads in §C.5, and the fact-processing and decision heads in §C.6, by performing causal intervention experiments in different sections of the input context.
  • For the queried-rule locating heads, we perform the following causal intervention experiments. In the original-altered prompt pairs, we only swap the location of the two rules in Rules section while keeping everything else the same. If these heads indeed perform their functions as we described, then when we run the model on the clean prompts, patching in the altered key activations at these heads (within the Rules section) should cause “negative” change to the model’s output, since it will cause these heads to mistake the queried-rule location in the altered prompt to be the right one, consequently storing the wrong rule information at the QUERY position. In particular, the model’s cross-entropy loss with respect to the original target should increase.
    Indeed, we observe in Figure 17 that interventions at these heads cause significantly larger increase in the cross-entropy loss than the other attention heads.
  • To further understand the fact-processing and decision heads, we construct original-altered pairs by only exchanging the truth value of the two nodes for the OR chain ([True, False]$\to$[False, True]). We perform subcomponent patching (query, key and value) of all attention heads in the Facts section, and found that out of all the patching results, only the fact-processing heads’ key activations exhibit strong “belief-altering” effects (large intervened logit difference). This indicates that, (1) the attention pattern of these heads are facts-sensitive, and (2) they perform (preliminary) filtering of the truth-value assignments, instead of blindly moving the truth-value assignments to the “:” position. These results are shown in Figure 18.
    Moreover, we also perform subcomponent patching at the “:” position, and found that only the decision heads’ query activations show nontrivial causal influence on the model’s correct output. This provides further evidence that these are the heads which relies on the embedding in the residual stream at the “:” position to send the correct answer tokens out. This result is shown in Figure 19.
5 Conclusion
[…]
Appendix
Appendix A Propositional logic problem and examples

### v4

@@ 4 Related Work
Coarse-Grained Evaluation of Reasoning in LLMs. There has been a flurry of work on thinking and reasoning models (Muennighoff et al., 2025; Ye et al., 2025; Wang et al., 2025). This extends the work on evaluating the reasoning abilities of LLMs across different tasks (Xue et al., 2024; Tafjord et al., 2021; Hendrycks et al., 2021; Chen et al., 2024; Patel et al., 2024; Berglund et al., 2024; Morishita et al., 2023; Fu et al., 2024; Seals & Shalin, 2024; Joren et al., 2025; Zhang et al., 2024, 2023; Saparov & He, 2023; Saparov et al., 2024; Sun et al., 2023; Luo et al., 2024; Arora et al., 2024; Han et al., 2024; Dziri et al., 2024; Yang et al., 2024). While these studies primarily benchmark their performance on sophisticated tasks, our work focuses on understanding “how” transformers reason on logic problems accessible to fine-grained analysis (e.g., mechanistic circuit analysis).
Fine-Grained Analysis of How LLMs Reason. There are fewer studies that provide fine-grained analysis of how LLMs reason. We build on ideas from mechanistic interpretability for arithmetic (Stolfo et al., 2023; Yu & Ananiadou, 2024; Kantamneni & Tegmark, 2025; Zhou et al., 2024), syllogistic reasoning (Kim et al., 2024), ontology CoT (Dutta et al., 2024), graph connectivity (Saparov et al., 2025), VLM tasks (Venhoff et al., 2025), and indirect object identification (Wang et al., 2023a). However, none studies problems with [Variable relationships]+[Variable value assignment]+[Query] aspects while considering pre-trained LLMs. We complement existing work on mechanistic analysis for such symbolic reasoning that considers small models trained on the task (Brinkmann et al., 2024; Guo et al., 2025b; Wang et al., 2024a; Zhu et al., 2024).
5 Discussion and Conclusion
@@
Apriori, it was not clear whether we could find isolated reasoning components in LLMs. In fact, we had a different conjecture based on studying the logic problem for smaller models, such as 3-layer transformers (results in Appendix D). For these small models, we found less isolation in their components, and in particular, the are not lazy in their processing. The second layer’s residual stream already contains a linearly decodable partial answer. This non-lazy behavior on Transformers trained specifically on a synthetic task is also supported by the work of Ye et al. (2024). Real-world LLMs seem to behave in a different manner, where not much processing happens until the QUERY token is seen. Fact processing and decisions depend on queried-rule locating and mover heads. Plainly, LLMs not only ingest facts from the input prompt (via in-context learning), but they also have internal components that process the provided information before downstream computation.
Limitations. Our circuit analysis is mostly at the level of attention patterns. It would be ideal to strengthen our findings with an analysis of the embedding level as well. Another area of improvement is comparing across problem complexity. With more computational resources, we should verify that the same components are integral when solving problem with longer logical chains. This would strengthen our claims about the reasoning circuits. Finally, there are unanswered quetions about how model size and family impact the types of reasoning circuits. We compare the results on Gemma with Mistral-7B in Appendix B.4.1, providing some evidence that model size may be more indicative of capabilities than model family (given certain similarities between Gemma-2-9B and Mistral-7B). Error analysis and steering are other important future directions, as is evaluating longer logical chains.
Appendix A Methodology and Problem Details
A.1 Appendix Table of Contents and Outline
In this section, we provide an overview of our methodology that we presented at a high level in Section 2. We then give forward pointers to our additional experimental evidence. We aim to provide an outline of the LLM analysis, guiding the reader through the rest of the appendix.
  • §A.2 gives representative examples of the propositional logic problem that we study. In general, we vary the variable names, the logical operator (“or” vs. “and”), the truth values of the variables (i.e., the “Facts”), and the QUERY variable.
  • §B.1 provides details of the few-shot examples and input format, and some initial results about model accuracies. In particular, we engineer the in-context examples to ensure that all models perform much better than random guessing.
  • §B.2 is an expository, background section that shows how activation patching can lead to circuit analysis. For completeness, we illustrate the main techniques on a toy 2-layer model, with figures. The main technique is to run the LLM on altered (alt) prompts, while freezing certain components to the activations on the original (orig) prompt. Throughout, we also define some notation, such as $\Delta_{orig\to alt}^{\mathcal{C}}$, which is the intervened logit difference. The relationship between the orig and alt prompts depends on the specific experiment, where we may change the query variable, the rule locations, or the truth values.
  • §B.3 defines the metrics for QUERY-based activation patching.
  • §B.4 through §B.7 provides the full analysis of the Mistral 7B model. We break this down into multiple subsections to present the details of the experiments, where we use the same approach for the Gemma models later on. In particular, §B.5 has the sufficient test, §B.6 contains results for varying the rule locations, and §B.7 has results for varying the facts.
  • §B.8 contains the experiments for Gemma 2 9B. This provides full details for the experiments that we report in the main paper.
  • §B.9 similarly contains results for Gemma 2 27B, including attention head analysis, and a discussion of the mechanical differences between Gemma 2 9B and 27B.
  • §C presents a full mechanistic analysis of small transformer models (e.g., 3-layer). We present the main insights in §C and then supplement this with the full details in §D. We train several models autoregressively on the propositional logic task. Then, we explore the mechanisms through a combination of experiments (patching, probing, classification, etc), focusing on a 3-layer model. We present the high level strategy of the 3-layer model in Figure 29, providing a contrast to the modular circuit in the pre-trained LLMs.
A.2 Propositional logic problem and examples